Find the equivalent capacitance.


C1=3 µF, C2=5 µF, C3=2 µF, C4=2 µF, C5=4 µF, C6=6 µF



1
Expert's answer
2021-03-29T09:09:43-0400

if the capacitors are in parallel, equivalent capacitance "=3+5+2+2+4+6=22" µF

if the capacitors are in series, "1\/C_E=1\/3+1\/5+1\/2+1\/2+1\/4+1\/6=39\/20"

"C_E =20\/39=0.5128\u00b5F"